# Segment Change Model (SCM) for Unsupervised Change detection in VHR Remote Sensing Images: a Case Study of Buildings

### v1

@@ 2.3 Piecewise Semantic Attention
building
non-building
roof, rooftop, building, house, apartment, residential, factory
baseball, diamond, bareland, swimming pool, basketball court, roundabout, playground
@@
$\begin{split}P_{bld}&=\sum\nolimits_{c\in\{bld\ related\ cls\}}p_{c}\\ \end{split}$ (2)
@@ 3 Experiments
F1
mIoU
OA
F1
mIoU
OA
PCA-KM[2]
39.22
28.92
54.13
41.85
31.73
59.04
CNN-CD[3]
45.49
35.41
64.78
43.13
33.70
63.20
DSFA[5]
47.65
40.70
77.33
48.39
42.46
81.02
DCVA[4]
52.89
46.09
84.75
53.90
44.89
79.44
GMCD[1]
50.02
42.75
79.64
55.55
47.56
84.49
SCM (Ours)
62.80 53.67 88.80 61.81 52.14 86.14
@@
F1
mIoU
OA
F1
mIoU
OA
base
42.21
30.12
53.35
42.24
30.61
54.87
base+RFF
52.59
41.80
72.40
52.17
42.13
74.16
SCM
62.80 53.67 88.80 61.81 52.14 86.14

### v2

@@ 2.3 Piecewise Semantic Attention
building non-building
roof, rooftop, building, house, apartment, residential, factory baseball, diamond, bareland, swimming pool, basketball court, roundabout, playground
@@
$\begin{split}P_{bld}&=\sum\nolimits_{c\in\{bld\ related\ cls\}}^{\ }p_{c}\\ \end{split}$ (2)
@@ 3 Experiments
F1 mIoU OA F1 mIoU OA
PCA-KM[2] 39.22 28.92 54.13 41.85 31.73 59.04
CNN-CD[3] 45.49 35.41 64.78 43.13 33.70 63.20
DSFA[5] 47.65 40.70 77.33 48.39 42.46 81.02
DCVA[4] 52.89 46.09 84.75 53.90 44.89 79.44
GMCD[1] 50.02 42.75 79.64 55.55 47.56 84.49
SCM (Ours) 62.80 53.67 88.80 61.81 52.14 86.14
@@
F1 mIoU OA F1 mIoU OA
base 42.21 30.12 53.35 42.24 30.61 54.87
base+RFF 52.59 41.80 72.40 52.17 42.13 74.16
SCM 62.80 53.67 88.80 61.81 52.14 86.14